A cliff diver from the top of a 100 m cliff. He begins his dive by jumping up with a velocity of 5 m/s. What is his velocity right before he hits the water?


a. 22.3 m/s

b. 11.15 m/s

c. 1,989 m/s

d. 44.6 m/s

 


1
Expert's answer
2022-02-15T09:58:35-0500

Given:

"v_1=5\:\rm m/s"

"h=\rm 100\: m"


The law of conservation of energy says

"\frac{mv_1^2}{2}+mgh=\frac{mv_2^2}{2}"

Thus, the final velocity

"v_2=\sqrt{v_1^2+2gh}=\sqrt{5^2+2*9.8*100}=\rm 44.6\: m/s"

Answer: d. 44.6 m/s